Consider two different 3-dimensional waves traveling through a section of the earth (e.g. froman earthquake). If the first wave has 5.0 times the energy of the second, what is the ratio ofamplitudes of the two waves?
Answer #1

As we know that -

Power (P) = Energy per second (E)

And the power is proportional to intensity.

Also, intensity is proportional to square of amplitude

This means -

Energy (E) is proportional to square of amplitude

So -

E1 / E2 = (A1 / A2)^2

=> A1 / A2 = sqrt(5) = 2.236

So, the ratio of amplitudes of the two waves = 2.236 (Answer)